in how many ways can 9 friends be formed into groups of 4 persons to play cards?
at first i thought it's just 9C4. then i realized another group of 4 can be made out of 5 left persons, leaving one behind. so the answer must be (9C4)(5C4) ways...? i really don't know, please help meeeh~

Answer by rothauserc(4718)

this is a combination problem because order does not matter, that is, set (1,2,3,4) is the same as set (2,3,4,1)
the total number of combinations can be calculated from the following formula
C = n! / (r! * (n - r)!) where n is the number of things to choose from and you choose r of them, therefore
C = 9! / (4! * (9 - 4)!
C = 9! / (4! * 5!)
C = 9*8*7*6 / 4*3*2*1 = 9*7*2 / 1 = 9*7*2 = 126 combinations